OpenStudy (anonymous):

Find the circumference of a circle whose area is 60 pi cm squared

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

no can you help

OpenStudy (skullpatrol):

You need to memorize two important formulas: $$\Huge A=\pi r^2$$ $$\Huge C =2 \pi r$$

OpenStudy (isaiah.feynman):

\[Area = \frac{ c^{2} }{ 4\pi } \] Where c is circumference.

OpenStudy (anonymous):

i dont see those answers there

OpenStudy (anonymous):

48.669=C

OpenStudy (skullpatrol):

A = pi*r^2 and C = 2*pi*r

OpenStudy (anonymous):

Kenny likes to count his money by tiling it on his table as shown. The diameter of a quarter is 24.26 mm. Find the area of the table not covered in quarters. Round your answer to the nearest hundredth. A. 3,031.28 mm2 B. 6,062.56 mm2 C. 12,295.98 mm2 D. 30,250.30 mm2

OpenStudy (anonymous):

what about this

OpenStudy (isaiah.feynman):

The circumference in terms of the area is given by \[c=2\pi \sqrt{\frac{ A }{ \pi }} \]Where A is area.

OpenStudy (anonymous):

Kenny likes to count his money by tiling it on his table as shown. The diameter of a quarter is 24.26 mm. Find the area of the table not covered in quarters. Round your answer to the nearest hundredth. A. 3,031.28 mm2 B. 6,062.56 mm2 C. 12,295.98 mm2 D. 30,250.30 mm2

OpenStudy (anonymous):

do you know this

OpenStudy (anonymous):

this I do not know.